# Designing LLM Chains by Adapting Techniques from Crowdsourcing Workflows

### v3

@@ 4.2.2. Subtasks
Subtasks are singular tasks in a workflow. Although highly varied, they follow several thematic functions: to generate, evaluate, transform, focus, and partition.
@@
Generate. Most workflows involve at least one step in which workers generate content. This generation may be writing words or code (Pavel et al., 2014; Weir et al., 2015; Zhang et al., 2022; Li et al., 2023; Zhang et al., 2023a) brainstorming and ideation (Chilton et al., 2019; Lasecki et al., 2013; Mahyar et al., 2018; Huang et al., 2018; Long et al., 2023; Yao et al., 2023), or answering a question or solving a problem (Kobayashi et al., 2018; Wu et al., 2023b; Xie et al., 2023; Eisenstein et al., 2022). Generation tasks may also ask for labels or annotations on given content (Guo et al., 2017, 2019; Chilton et al., 2014; Chang et al., 2017; Bragg et al., 2013), to translate between languages (Ambati et al., 2012; Zaidan and Callison-Burch, 2011), to contribute to a conversation with other workers (Mahyar et al., 2018; Wang et al., 2022; Liang et al., 2023; Du et al., 2023), to demonstrate a behavior (Bragg et al., 2021; Lee et al., 2017), or to source content from external sources (Kim et al., 2014; Chilton et al., 2019; Hahn et al., 2016; Kim and Monroy-Hernandez, 2016; Pietrowicz et al., 2013; Huang et al., 2015; Luther et al., 2015; Christoforaki and Ipeirotis, 2014). Some subtasks also require an explanation along with the generated content (Chang et al., 2017; Kim et al., 2017; Cheng and Bernstein, 2015; Drapeau et al., 2016; Schick et al., 2022; Gero et al., 2023; Fleming et al., 2023).
@@
Evaluate. One of the most common evaluation tasks is choosing the best among multiple possible options (Guo et al., 2017; Weir et al., 2015; Guo et al., 2019; Cheng and Bernstein, 2015; Kittur et al., 2012; Liu et al., 2018). Other subtasks may include rating and ranking among multiple options (Pavel et al., 2014; Salehi et al., 2017; Marcus et al., 2011a; Huffaker et al., 2020; Willett et al., 2012; Mahyar et al., 2018) or a pairwise comparison (Mohanty et al., 2019; Little et al., 2010b, a; Marcus et al., 2011a, b; Parameswaran et al., 2023; Bursztyn et al., 2022). Evaluation subtasks may also determine if another round of generation needs to occur (Wu et al., 2022b; LaToza et al., 2014; Wang et al., 2022; Liang et al., 2023).
@@
Workflows can decide what moves forward through majority vote (Huang et al., 2017; Dai et al., 2010; Lasecki et al., 2015; Liu et al., 2018; Kim et al., 2014; Guo et al., 2019), other aggregations of voting tasks (e.g., expectation maximization (Drapeau et al., 2016; Parameswaran et al., 2023), weighted mean (Mohanty et al., 2019), weak supervision (Arora et al., 2022)), returning the highest ranked or rated items (Pavel et al., 2014; Huffaker et al., 2020), or automated methods. These automated methods include measuring agreement from the generation step (Liu et al., 2018; Hahn et al., 2016; Huffaker et al., 2020; Chang et al., 2017; Huang et al., 2015; Nair et al., 2023b; Bernstein et al., 2011) and using non-AI checks like semantic embeddings and rule-based heuristics (Parameswaran et al., 2023; Wang et al., 2023; Ambati et al., 2012). There are also subtasks that seek critiques and reviews from the user (Nebeling et al., 2016; Schick et al., 2022) or another worker (Kim et al., 2017; Kulkarni et al., 2014; Valentine et al., 2017; Christoforaki and Ipeirotis, 2014; Mahyar et al., 2018; Schick et al., 2022; Parameswaran et al., 2023; Madaan et al., 2023b; Liang et al., 2023). These subtasks may ask for feedback or verify if errors or other qualities exist (Huang et al., 2023b; Bragg et al., 2013; Christoforaki and Ipeirotis, 2014; Chilton et al., 2019; Weir et al., 2015; Huang et al., 2015; Kim et al., 2014; Gero et al., 2023; Huang et al., 2017; Bernstein et al., 2010; Yang et al., 2022; Wu et al., 2023c; Kittur et al., 2011; Liang et al., 2023; Kulkarni et al., 2012; André et al., 2013; Butler et al., 2017).
@@
Transform. Other subtasks transform given data by merging or improving it. Some subtasks merge multiple streams of incoming content. This merge could be algorithmic (Alshaibani et al., 2020; Ahmad et al., 2011; Alshaibani and Quinn, 2021; Teevan et al., 2016; Chilton et al., 2019; Kittur et al., 2011) or accomplished by the worker (Kittur et al., 2011; Wu et al., 2023c, 2022b; Lee et al., 2017; LaToza et al., 2014; Ambati et al., 2012; Teevan et al., 2016; Kittur et al., 2012; Wu et al., 2023c; Pietrowicz et al., 2013; Luther et al., 2015; Kulkarni et al., 2012). Other subtasks improve content generated by previous steps, sometimes by incorporating feedback from evaluation steps. Improving on previously generated content can come in many forms: updating the format of the content (Huang et al., 2023b; Wu et al., 2022b; Eisenstein et al., 2022; Wang et al., 2022; Arora et al., 2022), pruning bad results (Liu et al., 2018; Gero et al., 2023), adding missing results (Huang et al., 2023b; Liu et al., 2018; Weir et al., 2015), correcting errors (Huang et al., 2023b; Yang et al., 2022; Drapeau et al., 2016; Kobayashi et al., 2018; Gouravajhala et al., 2018), and performing direct improvements (Madaan et al., 2023b; Hahn et al., 2016; Mahyar et al., 2018; Lee et al., 2017; Huffaker et al., 2020; Nebeling et al., 2016; Christoforaki and Ipeirotis, 2014; Little et al., 2010b; Liem et al., 2011; Dai et al., 2010; Schick et al., 2022; Kim et al., 2017; Bernstein et al., 2010; Guo et al., 2019; McDonnell et al., 2016).
@@
Focus. One set of subtasks looks to select the attention of the upcoming tasks within a broader context. For example, this step may find patches of the input that need editing, extract information from context, or focus on common keywords (Bernstein et al., 2011; Huffaker et al., 2020; Gero et al., 2023; Creswell and Shanahan, 2022; Creswell et al., 2022; Nair et al., 2023b; Kim et al., 2017).
@@
Partition. Some subtasks transform a task by breaking it up into smaller pieces (Kittur et al., 2011; LaToza et al., 2014; Kulkarni et al., 2012; Wu et al., 2023c; Zhou et al., 2022b; Valentine et al., 2017; Wu et al., 2022b; Huang et al., 2023c). For example, the Divide step in the Turkomatic workflow asks workers to divide a given subtask into simpler subtasks. The workflow calls this step recursively until the subtask complexity is valued (by a separate crowdworker) at a specified cost (Kulkarni et al., 2012).
Three columns headed by ”If…”, ”Then…” and ”With…”. The ”If” column contains different possible challenges relating to achieving objectives. The ”Then” column lists all of the strategies. There are arrows connecting the ”If” column to the ”Then” column, with arrows from ”navigating a large creative space” objectives challenge to the adaptive architectures, diverse responses, and user guidance strategies. The ”With” column lists all of the tactics. There are arrows connecting the ”Then” column to the ”With” column. There are many arrows, and they correspond to the tactics used for strategies as defined in the text and other figures.
@@ 4.2.3. Workflow architecture
Sequential architectures order subtasks in a chain in which the output of one worker passes forward to another worker (Wang et al., 2023; Nair et al., 2023a; Wu et al., 2022b; Huang et al., 2023a; Zhou et al., 2022b; Huang et al., 2023b; Reppert et al., 2023; Gero et al., 2023; Wu et al., 2023c; Long et al., 2023; Eisenstein et al., 2022).
@@
Branching workflows apply conditionals to choose different execution paths based on the outputs of prior steps (Huang et al., 2018, 2023b; Kittur et al., 2012; Wu et al., 2022a; Zharikova et al., 2023).
@@
Redundant architectures require different workers to complete the same task in parallel or iteratively. In parallel generation, multiple workers perform the same subtask separately (Weir et al., 2015; Guo et al., 2019; Cobbe et al., 2021; Butler et al., 2017; Lin et al., 2014a; Alshaibani and Quinn, 2021; Huang et al., 2017). This subtask is often, but not always, followed by an aggregation or evaluation subtask to merge or select among the different parallel generations. Iterative architectures run iteratively over a changing input. For example, iterative improvement takes the output of a subtask and iteratively gives it to a series of workers to improve, sometimes with interspersed decision tasks choosing between the current and prior iterations (Luther et al., 2015; André et al., 2014; LaToza et al., 2014; Goto et al., 2016; Dai et al., 2013; Lasecki et al., 2015; Willett et al., 2012). The number of iterations can be determined to be a fixed number of steps (Wu et al., 2023c; Kim et al., 2017; Huffaker et al., 2020; Little et al., 2010b; Zaidan and Callison-Burch, 2011), by including a subtask to check for sufficiency (Madaan et al., 2023b; Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b; Wang et al., 2022; Wu et al., 2023b; Liang et al., 2023; Salehi et al., 2017; Kulkarni et al., 2014; Guo et al., 2016), until the user stops calling iterations (Cheng and Bernstein, 2015; Reppert et al., 2023; Kulkarni et al., 2014; Chilton et al., 2019; Willett et al., 2012), until the outputs converge (Schick et al., 2022; Liem et al., 2011; Du et al., 2023; Gero et al., 2023; Liem et al., 2011), until all quality checks are passed (Kim et al., 2017; Bragg et al., 2013; Chilton et al., 2013; Marcus et al., 2011a, b; Yang et al., 2022; Teevan et al., 2016), from increasing complexity levels (Alshaibani and Quinn, 2021; Alshaibani et al., 2020; Bursztyn et al., 2022), and through decision-theoretic approaches (Park et al., 2022; Dai et al., 2010).
@@
Dynamic architectures support changing the workflow architecture as it is executed. Map-reduce is a dynamic architecture in which workers break down a task recursively until the subtasks are an appropriate size (Wu et al., 2023c; Liem et al., 2011; Kittur et al., 2011; Kim and Monroy-Hernandez, 2016; Kulkarni et al., 2012; Huang et al., 2023c; Kaur et al., 2017; Valentine et al., 2017; Liem et al., 2011). The literature defines such workflows as a hierarchy (Huang et al., 2023b; Kittur et al., 2011). Other dynamic workflow architectures include doing beam search across multiple options (Xie et al., 2023; Yao et al., 2023; Creswell and Shanahan, 2022) and allowing for workers to flexibly create tasks on a shared to-do list before choosing which of those to complete (Chilton et al., 2019; Nebeling et al., 2016; Mahyar et al., 2018; Zhang et al., 2012; Kittur et al., 2011; LaToza et al., 2014). Some other architecture methods may have dynamic elements, such as parallelizing an input into chunks and running the workflow for each chunk (André et al., 2013; Gouravajhala et al., 2018; Zaidan and Callison-Burch, 2011; Kim et al., 2014; Pavel et al., 2014; Kaur et al., 2017; Bernstein et al., 2010) or dynamically determining the number of iterative steps (Chen et al., 2019; Chilton et al., 2014; Madaan et al., 2023b; Zhou et al., 2022b; Nair et al., 2023b).

### v4

@@ 4.2.2. Subtasks
Subtasks are singular tasks in a workflow. Although highly varied, they follow several thematic functions: to generate, evaluate, improve, focus, partition, and merge.
@@
Generate. Most workflows involve at least one step in which actors generate content. This generation may be writing words or code [C(Weir et al., 2015; Pavel et al., 2014) L(Li et al., 2023; Zhang et al., 2022, 2023a)] brainstorming and ideating [C(Chilton et al., 2019; Lasecki et al., 2013; Huang et al., 2018; Mahyar et al., 2018) L(Yao et al., 2024; Long et al., 2023)], or answering a question or solving a problem [C(Kobayashi et al., 2018) L(Xie et al., 2023; Eisenstein et al., 2022; Wu et al., 2024)]. Generation tasks may also ask for labels or annotations on given content [C(Chilton et al., 2014; Guo et al., 2019; Bragg et al., 2013; Chang et al., 2017; Guo et al., 2017)], to translate between languages [C(Ambati et al., 2012; Zaidan and Callison-Burch, 2011)], to contribute to a conversation with other actors [C(Mahyar et al., 2018) L(Liang et al., 2023; Du et al., 2023; Wang et al., 2022)], to demonstrate a behavior [C(Bragg et al., 2021; Lee et al., 2017)], or to source content from external sources [C(Luther et al., 2015; Pietrowicz et al., 2013; Chilton et al., 2019; Kim et al., 2014; Christoforaki and Ipeirotis, 2014; Kim and Monroy-Hernandez, 2016; Hahn et al., 2016; Huang et al., 2015)]. Some subtasks also require an explanation along with the generated content [C(Kim et al., 2017; Cheng and Bernstein, 2015; Drapeau et al., 2016; Chang et al., 2017) L(Gero et al., 2023; Schick et al., 2022; Fleming et al., 2023)].
@@
Evaluate. One of the most common evaluation tasks is choosing the best among multiple possible options [C(Liu et al., 2018; Weir et al., 2015; Guo et al., 2019; Kittur et al., 2012; Cheng and Bernstein, 2015; Guo et al., 2017) L(Mirowski et al., 2023; Yang et al., 2022; Yao et al., 2024)]. Other subtasks may include rating and ranking among multiple options [C(Marcus et al., 2011a; Pavel et al., 2014; Willett et al., 2012; Salehi et al., 2017; Mahyar et al., 2018; Huffaker et al., 2020) L(Yao et al., 2024; Li et al., 2023; Parameswaran et al., 2023)] or a pairwise comparison [C(Little et al., 2010a; Marcus et al., 2011a; Little et al., 2010b; Marcus et al., 2011b; Mohanty et al., 2019) L(Bursztyn et al., 2022; Parameswaran et al., 2023)]. Evaluation subtasks may also determine if another round of generation needs to occur [C(LaToza et al., 2014) L(Liang et al., 2023; Wang et al., 2022; Wu et al., 2021)].
@@
Workflows can decide what moves forward through majority vote [C(Liu et al., 2018; Lasecki et al., 2015; Kim et al., 2014; Guo et al., 2019; Huang et al., 2017; Dai et al., 2010)], other aggregations of voting tasks (e.g., expectation maximization [C(Drapeau et al., 2016) L(Parameswaran et al., 2023)] or weighted mean [C(Mohanty et al., 2019)]), returning the highest ranked or rated items [C(Pavel et al., 2014; Huffaker et al., 2020)], or automated methods. These automated methods include measuring agreement from another step [C(Liu et al., 2018; Hahn et al., 2016; Chang et al., 2017; Huang et al., 2015; Bernstein et al., 2011; Huffaker et al., 2020) L(Nair et al., 2023b)], weak supervision[L(Arora et al., 2023)], and using non-AI checks like semantic embeddings and rule-based heuristics [C(Ambati et al., 2012) L(Parameswaran et al., 2023; Wang et al., 2023)]. There are also subtasks that seek critiques and reviews from the user [C(Nebeling et al., 2016) L(Schick et al., 2022)] or a worker [C(Kulkarni et al., 2014; Valentine et al., 2017; Kim et al., 2017; Christoforaki and Ipeirotis, 2014; Mahyar et al., 2018) L(Parameswaran et al., 2023; Madaan et al., 2024; Liang et al., 2023; Schick et al., 2022)]. These subtasks may ask for feedback or verify if errors or other qualities exist [C(Chilton et al., 2019; Weir et al., 2015; Kim et al., 2014; Christoforaki and Ipeirotis, 2014; André et al., 2013; Bragg et al., 2013; Huang et al., 2017; Bernstein et al., 2010; Huang et al., 2015; Kulkarni et al., 2012; Butler et al., 2017; Kittur et al., 2011) L(Huang et al., 2023b; Yang et al., 2022; Gero et al., 2023; Liang et al., 2023; Wu et al., 2023b)].
@@
Improve. Other subtasks improve content generated by previous steps, sometimes by incorporating feedback from evaluation steps. Improving on previously generated content can come in many forms: updating the format of the content [L(Huang et al., 2023b; Wang et al., 2022; Wu et al., 2021; Eisenstein et al., 2022; Arora et al., 2023)], pruning bad results [C(Liu et al., 2018) L(Gero et al., 2023)], adding missing results [C(Liu et al., 2018; Weir et al., 2015) L(Huang et al., 2023b)], correcting errors [C(Drapeau et al., 2016; Gouravajhala et al., 2018; Kobayashi et al., 2018) L(Huang et al., 2023b; Yang et al., 2022)], and performing direct improvements [C(Liem et al., 2011; Kim et al., 2017; Little et al., 2010b; Nebeling et al., 2016; Guo et al., 2019; McDonnell et al., 2016; Christoforaki and Ipeirotis, 2014; Hahn et al., 2016; Bernstein et al., 2010; Mahyar et al., 2018; Lee et al., 2017; Dai et al., 2010; Huffaker et al., 2020) L(Madaan et al., 2024; Schick et al., 2022)].
@@
Focus. One set of subtasks looks to select the attention of the upcoming tasks within a broader context. For example, this step may find patches of the input that need editing, extract information from context, or focus on common keywords [C(Kim et al., 2017; Bernstein et al., 2011; Huffaker et al., 2020) L(Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b; Gero et al., 2023)].
@@
Partition. Some subtasks transform a task by breaking it up into smaller pieces [C(Valentine et al., 2017; LaToza et al., 2014; Kulkarni et al., 2012; Kittur et al., 2011) L(Zhou et al., 2022b; Wu et al., 2021, 2023b; Huang et al., 2023c)]. For example, the Divide step in the Turkomatic workflow asks crowdworkers to divide a given subtask into simpler subtasks. The workflow calls this step recursively until the subtask complexity is valued (by a separate crowdworker) at a specified cost [C(Kulkarni et al., 2012)].
Merge. Subtasks can merge multiple streams of incoming content. This merge could be algorithmic [C(Chilton et al., 2019; Alshaibani et al., 2020; Teevan et al., 2016; Alshaibani and Quinn, 2021; Kittur et al., 2011; Ahmad et al., 2011)] or accomplished by an actor [C(Luther et al., 2015; Pietrowicz et al., 2013; Ambati et al., 2012; LaToza et al., 2014; Teevan et al., 2016; Kittur et al., 2012; Kulkarni et al., 2012; Lee et al., 2017; Kittur et al., 2011) L(Wu et al., 2021, 2023b)].
Two sets of bar graphs. The left is titled ”Objectives Distribution” and contains data on the objectives. The right is titled ”Tactics Distribution” and contains data on the tactics. The x-axis of each contains the design space categories, and the y-axis ranges from 0 percent to 100 percent. Each design space category on the x axis has a bar color-coded to represent the crowdsourcing or LLM literature. All categories, besides the “LLM” and “Crowdworker” categories, exist in both literatures. The bar values summarize the categorizations found in the supplementary.
@@ 4.2.3. Workflow architecture
Sequential architectures order subtasks in a chain in which the output of one actor passes forward to another actor [C(Christoforaki and Ipeirotis, 2014; McDonnell et al., 2016; Huang et al., 2017; Huffaker et al., 2020) L(Zhou et al., 2022b; Huang et al., 2023a, b; Nair et al., 2023a; Wu et al., 2021; Eisenstein et al., 2022; Gero et al., 2023; Long et al., 2023; Reppert et al., 2023; Wu et al., 2023b; Wang et al., 2023)].
@@
Branching workflows apply conditionals to choose different execution paths based on the outputs of prior steps [C(Huang et al., 2018; Kittur et al., 2012) L(Huang et al., 2023b; Wu et al., 2022a; Zharikova et al., 2023)].
@@
Redundant architectures require different actors to complete the same task in parallel or iteratively. In parallel generation, multiple actors perform the same subtask separately [C(Weir et al., 2015; Guo et al., 2019; Huang et al., 2017; Butler et al., 2017; Alshaibani and Quinn, 2021; Lin et al., 2014a) L(Cobbe et al., 2021)]. This subtask is often, but not always, followed by an aggregation or evaluation subtask to merge or select among the different parallel generations. Unlike parallelization that runs a similar subtask on different items at the same time (e.g., five crowdworkers label one image each for five labeled images total), parallel redundant architectures run the same subtask on the same item (e.g., five crowdworkers independently label the same image for one image total). Iterative architectures run iteratively over a changing input. For example, iterative improvement takes the output of a subtask and iteratively gives it to a series of actors to improve, sometimes with interspersed decision tasks choosing between the current and prior iterations [C(Luther et al., 2015; Lasecki et al., 2015; André et al., 2014; LaToza et al., 2014; Goto et al., 2016; Willett et al., 2012; Dai et al., 2013)]. The number of iterations can be determined to be a fixed number of steps [C(Kim et al., 2017; Little et al., 2010b; Zaidan and Callison-Burch, 2011; Huffaker et al., 2020) L(Wu et al., 2023b)], by including a subtask to check for sufficiency [C(Kulkarni et al., 2014; Salehi et al., 2017; Guo et al., 2016) L(Creswell et al., 2022; Wang et al., 2022; Creswell and Shanahan, 2022; Wu et al., 2024; Nair et al., 2023b; Liang et al., 2023; Madaan et al., 2024)], until the user stops calling iterations [C(Kulkarni et al., 2014; Chilton et al., 2019; Cheng and Bernstein, 2015; Willett et al., 2012) L(Reppert et al., 2023)], until the outputs converge [C(Liem et al., 2011) L(Schick et al., 2022; Gero et al., 2023; Du et al., 2023)], until all quality checks are passed [C(Kim et al., 2017; Marcus et al., 2011a, b; Bragg et al., 2013; Chilton et al., 2013; Teevan et al., 2016) L(Yang et al., 2022)], through increasing complexity levels [C(Alshaibani et al., 2020; Alshaibani and Quinn, 2021)], and through decision-theoretic approaches [C(Dai et al., 2010) L(Park et al., 2022)].
@@
Dynamic architectures support changing the workflow architecture as it is executed. Map-reduce is a dynamic architecture in which actors break down a task recursively until the subtasks are an appropriate size [C(Liem et al., 2011; Valentine et al., 2017; Kaur et al., 2017; Kim and Monroy-Hernandez, 2016; Kulkarni et al., 2012; Kittur et al., 2011) L(Wu et al., 2023b; Huang et al., 2023c)]. The literature defines such workflows as a hierarchy [C(Kittur et al., 2011)]. Other dynamic workflow architectures include doing beam search across multiple options [L(Yao et al., 2024; Creswell and Shanahan, 2022; Xie et al., 2023)] and allowing for actors to flexibly create tasks on a shared to-do list before choosing which of those to complete [C(Chilton et al., 2019; Nebeling et al., 2016; LaToza et al., 2014; Mahyar et al., 2018; Kittur et al., 2011; Zhang et al., 2012)]. Some other architecture methods may have dynamic elements, such as parallelizing an input into chunks and running the workflow for each chunk [C(Kaur et al., 2017; Kim et al., 2014; André et al., 2013; Pavel et al., 2014; Zaidan and Callison-Burch, 2011; Gouravajhala et al., 2018)] or dynamically determining the number of iterative steps [C(Chilton et al., 2014; Chen et al., 2019) L(Zhou et al., 2022b; Nair et al., 2023b; Madaan et al., 2024)].